243
edits
ElenaZhivun (talk | contribs) |
ElenaZhivun (talk | contribs) |
||
We observed response in both X and Y channels of the lock-in. The X response remained positive, and changed by 20-50%. The Y response was an 10 times smaller than the X response, and changed sign as the optical frequency changed. The pump is parallel to the Z direction and orthogonal to the probe. It appears that the measured magnetic field noise has contributions both from the changes in atomic polarization and light shifts, but the light shifts contributions are an order of magnitude smaller.
== 11/10/2016 ==
=== Optimizing magnetometer performance in WIMR ===
Tuned the probe and pump lasers to get the optimal magnetometer signal. The magnetometer appears to have residual sensitivity to the magnetic field when the pump is turned off. After tuning the probe further away from the resonance and decreasing the probe power, we could get nearly flat noise at 1 fT/sqrt(Hz) in Y mode.
==== Before (large probe power, close to the optical transition ====
TODO: add the plot
==== After (reduced probe power, detuned from the optical transition ====
TODO: add the plot
=== Popcorn noise in the diff/sum circuit ===
There are strange noise peaks in the time series in the probe noise. They appear only in the diff/sum mode of the polarimeter circuit, and go away if only the differential channel is used. They might be caused by a large digital gain (30), that results in the significant bits being lost and digitization noise appearing and passing through the filters.
==== Difference only ====
TODO: add the plot
==== Diff/sum ====
TODO: add the plot
=== fMCG live preview ===
Tested out the fMCG live preview in the FPGA magnetometer. The filter cuts out 60 Hz and 120 Hz and they are not visible in the preview. Then a large (almost of the size of the full dynamic range) 60 Hz disturbance is applied to the magnetic field, the filter eliminates it completely. For such a large disturbance, it takes several seconds until the 60 Hz exponentially disappears. The filter doesn't seem to distort the MCG signal significantly. The signal to noise on the MCG signal is poor, and the DC offset makes it hard to preview. I have replaced the 100 Hz low-pass filter with a 1-80 Hz band-pass filter to fix this problem, and make the preview signal match the pre-processed data in Saki as much as possible.
=== Digital current driver ===
Preliminary tests of the digital current driver circuit. Verified that the circuit works, but the components that Mike has found in Billy's can only run at 2 MHz (we need 20 or 40 MHz). We have also tried applying the digital modulation to the input of the existing current driver, and it works as expected. The atoms react to the applied field change and not to the the high-frequency noise. It is unclear whether the existing current driver is low-passing the digital signal, or the atoms.
=== Pump current stabilization ===
Wanted to check whether some part of the measured magnetic noise might actually be coming from the pump power noise. First we tried to stabilize the power by feeding back on the LD current, but that didn't work. The laser current was changing, but we didn't observe significant change of laser power after the tapered amplifier. The we locked the power on the tapered amplifier itself, and that didn't reduce the measured magnetic field noise. So either the room is too noisy, or the power stabilization wasn't enough.
|